. determine the mass of water that absorbs 1234 J of heat and undergoes a temperature change of 33.4oC.
1
Expert's answer
2013-01-21T07:25:45-0500
Heat is evaluated asQ=c*m*(t2-t1)
where m is mass and c is specific heat capacity.
In our case t2-t1=33.4. so we find mass
m = Q/(c*(t2-t1)) = 1234/(4200*33.4) = 0.0088 kg = 8.8g
mass of water is 8.8 g